Repeat algebraically the area of a rectangle with length= 2x+1 and width= x+3
Answer by nerdybill(7384)

length= 2x+1 and width= x+3
.
since area of any rectangle is width time length:
Area = (x+3)(2x+1)
Apply FOIL:
Area = 2x^2 + x + 6x + 3
Area = 2x^2 + 7x + 3